Diffraction grating A has 16 000 lines/cm, and diffraction grating B has 12 000 lines/cm. Light of the same wavelength is sent through both gratings to a screen located a distance L from the gratings. The spacing of the maxima for grating A is 3.46 cm. What is the spacing for the maximum for grating B?
1
Expert's answer
2020-06-17T09:34:47-0400

Given:

"d_A=1/16 000 \text{ cm}^{-1}\\
d_B=1/12 000 \text{ cm}^{-1}\\
x_A=3.46\text{ cm}."

Solution

The fringe spacing can be expressed as


"x=\frac{\lambda L}{d},"

where L - distance from the screen.

Thus, for A we have


"x_A=\frac{\lambda L}{d_A},\\
\lambda L=x_Ad_A."

For B:


"x_B=\frac{\lambda L}{d_B}=\frac{\lambda L}{d_B}=x_A\frac{d_A}{d_B}=3.46\cdot\frac{12000}{16000}=2.60\text{ cm}."